OpenStudy (anonymous):

∫[(e^-x-e^x)/(e^2x+e^-2x+2)] by substitution

OpenStudy (anonymous):

Hah! I see you trying to get one more calculus question in myin. ;p

myininaya (myininaya):

lol you too

OpenStudy (anonymous):

true.

OpenStudy (anonymous):

god I'm going to miss learning math

myininaya (myininaya):

i think we might have to do some algebra for doing a substitution let me see

OpenStudy (anonymous):

part of this looks kinda like a sinh(x)

OpenStudy (anonymous):

Or rather exactly like it.

myininaya (myininaya):

right i was thinking of but i forgot that part of calculus (or trig)

OpenStudy (anonymous):

\[sinh(x) = \frac{e^x - e^{-x}}{2}\]

OpenStudy (anonymous):

Although it's actually easier than that. The numerator is just the derivative of the denominator (missing a factor of 2)

OpenStudy (anonymous):

so just use that denominator for the substitution and I think it'll work out. too tired to do it fast enough ;p

myininaya (myininaya):

wait i think i might see something

myininaya (myininaya):

\[\int\limits_{}^{}\frac{e^{-2x}-1}{e^{2x}+e^{-2x}+2} e^x dx\] let u=e^x => du=e^x dx so u^{2}=e^{2x} so u^{-2}=e^{-2x} so we have \[\int\limits_{}^{}\frac{u^{-2}-1}{u^2-u^{-2}+2} du\]

myininaya (myininaya):

now let me see if i can work with this

myininaya (myininaya):

yep

myininaya (myininaya):

\[\int\limits_{}^{}\frac{u^2}{u^2}\frac{u^{-2}-1}{u^2-u^{-2}+2} du\]

myininaya (myininaya):

\[\int\limits_{}^{}\frac{1-u^2}{u^4-1+2u^2}du\]

myininaya (myininaya):

now we may be able to use partial fractions lets see if we can factor the denominator

myininaya (myininaya):

no we may have to use another subsitution

OpenStudy (anonymous):

Yeah, my way didn't work either.

myininaya (myininaya):

\[-\int\limits_{}^{}\frac{u^2-1}{u^4+2u^2-1}du\] i'm going to try a trig substituion

myininaya (myininaya):

\[-\int\limits_{}^{}\frac{u^2-1}{u^4+2u^2+1-2}du=-\int\limits_{}^{}\frac{u^2-1}{(u^2+1)^2-2}du\]

myininaya (myininaya):

\[\sec(\theta)=\frac{u^2+1}{\sqrt{2}}\]

myininaya (myininaya):

\[\sqrt{2} \sec(\theta)=u^2+1\] \[\sqrt{2} \sec(\theta)\tan(\theta) d \theta=2u du\]

myininaya (myininaya):

\[-\int\limits_{}^{}\frac{\sqrt{2}\sec(\theta)-2}{(\sqrt{2}\sec(\theta))^2-2} du\]

myininaya (myininaya):

\[-\int\limits\limits_{}^{}\frac{\sqrt{2}\sec(\theta)-2}{(\sqrt{2}\sec(\theta))^2-2} \sqrt{2} \sec(\theta) \tan(\theta) d \theta\]

myininaya (myininaya):

\[-\int\limits_{}^{}\frac{\sqrt{2} \sec(\theta)-2}{2(\sec^2(\theta)-1)} \sqrt{2} \sec(\theta) \tan(\theta) d \theta\]

myininaya (myininaya):

\[-\int\limits_{}^{}\frac{\sqrt{2}\sec(\theta)-2}{2\tan^2(\theta)} \sqrt{2}\sec(\theta)\tan(\theta) d \theta\]

myininaya (myininaya):

\[-\frac{\sqrt{2}}{2}\int\limits_{}^{}\frac{\sqrt{2}\sec(\theta)-2}{\tan(\theta)}\sec(\theta) d \theta\]

myininaya (myininaya):

\[-\frac{\sqrt{2}}{2}\int\limits_{}^{}\frac{\sqrt{2}\sec^2(\theta)-2\sec(\theta)}{\tan(\theta)} d \theta \]

myininaya (myininaya):

lol ok goodnight

OpenStudy (anonymous):

<3

OpenStudy (anonymous):

I just couldn't recall the derivatives off the top of my head.

OpenStudy (anonymous):

Had to look em up.

myininaya (myininaya):

i wonder if i was going a long way or a deadend way

OpenStudy (anonymous):

I dunno.

myininaya (myininaya):

wait i can separate that fraction

myininaya (myininaya):

\[-\frac{\sqrt{2}}{2}\int\limits\limits_{}^{}\frac{\sqrt{2}\sec^2(\theta)-2\sec(\theta)}{\tan(\theta)} d \theta \] \[=\frac{-\sqrt{2}}{2}(\sqrt{2}\int\limits_{}^{}\frac{\sec^2(\theta)}{\tan(\theta)} d \theta-2\int\limits_{}^{}\frac{\sec(\theta)}{\tan(\theta)}d \theta)\]

myininaya (myininaya):

\[y=\tan(\theta)=> dy=\sec^2(\theta) d \theta\]

OpenStudy (anonymous):

Hrm.. now I'm sure my answer is right. But I'm too tired to look at it now. I'll come back later.

myininaya (myininaya):

so looking at \[\int\limits_{}^{}\frac{\sec^2(\theta)}{\tan(\theta)}d \theta=\int\limits_{}^{}\frac{du}{u}=\ln|u|+C=\ln|\tan(\theta)|+C\]

OpenStudy (anonymous):

btw guys the answer is 1/(e^x+e^-x)+c

myininaya (myininaya):

\[\int\limits_{}^{}\frac{\sec(\theta)}{\tan(\theta)} d \theta=\int\limits_{}^{}\frac{\sin(\theta)}{\cos^2(\theta)} d \theta\] \[h=\cos(\theta)=> dh=-\sin(\theta) d \theta\] \[-\int\limits_{}^{}\frac{dh}{h^2}=-\frac{h^{-1}}{-1}+K=\frac{1}{h}+K=\frac{1}{\cos(\theta)}+K\]

myininaya (myininaya):

\[\frac{-\sqrt{2}}{2}(\sqrt{2}(\ln|\tan(\theta)|)-2(\frac{1}{\cos(\theta)}))+constant \] almost there

myininaya (myininaya):

remember \[\sec(\theta)=\frac{u^2+1}{\sqrt{2}} \] \[\cos(\theta)=\frac{\sqrt{2}}{u^2+1}\] \[\sin(\theta)=\frac{\sqrt{(u^2+1)^2-2}}{u^2+1}\] \[\tan(\theta)=\frac{\sqrt{(u^2+1)^2-2}}{\sqrt{2}}\]

myininaya (myininaya):

so now we have \[\frac{-\sqrt{2}}{2}(\sqrt{2}\ln|\sqrt{\frac{(u^2+1)^2-2}{2}}|-2\frac{u^2+1}{\sqrt{2}})\]

myininaya (myininaya):

ok but u=e^{x}

myininaya (myininaya):

\[\frac{-\sqrt{2}}{2}(\sqrt{2}*\frac{1}{2}\ln|\frac{(e^{2x}+1)^2-1}{2}|-\frac{2}{\sqrt{2}}(e^{2x}+1))+c\]

myininaya (myininaya):

now we need to simplify this

myininaya (myininaya):

\[-\frac{2}{4}\ln|\frac{(e^{2x}+1)^2-1}{2}|+e^{2x}+1+c\]

myininaya (myininaya):

\[\frac{-1}{2}(\ln|e^{4x}+2e^2x+1-1|-\ln(2))+e^{2x}+1+c\]

myininaya (myininaya):

\[\frac{-1}{2}\ln(e^{4x}+2e^{2x})+\frac{\ln(2)}{2}+e^{2x}+1+c\]

myininaya (myininaya):

\[\frac{-1}{2}\ln(e^{2x}(e^{2x}+2))+\frac{1}{2}\ln(2)+e^{2x}+1+c\] \[\frac{-1}{2}(lne^{2x}+\ln(e^{2x}+2))+\frac{1}{2}\ln(2)+e^{2x}+1+c\] \[-\frac{1}{2} \cdot 2 -\frac{1}{2}\ln(e^{2x}+2)+e^{2x}+1+c\]

myininaya (myininaya):

\[-\frac{1}{2}\ln(e^{2x}+2)+e^{2x}+c\]

OpenStudy (anonymous):

Why are you still here! ;p

myininaya (myininaya):

lol

myininaya (myininaya):

i seen how to finish mine way lol

myininaya (myininaya):

wonder if i check this what would i get back

myininaya (myininaya):

-exp(2*x)/(exp(2*x)+2)+2*exp(2*x) this is what i get back i wonder if i can rewrite it as what we started out with

myininaya (myininaya):

\[\frac{-e^{2x}}{e^{2x}+2}+2e^{2x}=\frac{-e^{2x}+2e^{2x}(e^{2x}+2)}{e^{2x}+2}\]

myininaya (myininaya):

\[=\frac{2e^{4x}+4e^{2x}-e^{2x}}{e^{2x}+1}=\frac{2e^{4x}+3e^{2x}}{e^{2x}+1}\] :(

OpenStudy (anonymous):

Hrm my answer is not right either.

myininaya (myininaya):

why would i make that 1 a 2 lol

myininaya (myininaya):

i probably made a mistake somewhere due to all the substitutions

myininaya (myininaya):

the more lenthiness=probability of mistakes increases

OpenStudy (anonymous):

getting out the whiteboard...

myininaya (myininaya):

ok i'm gonna do this on paper

OpenStudy (anonymous):

Ok, so found 2 mistakes so far in m original

myininaya (myininaya):

i found some mistakes in mine for sure lol

OpenStudy (anonymous):

Ah ha!

myininaya (myininaya):

i keep making mistakes

myininaya (myininaya):

i think my brain doesn't work anymore

OpenStudy (anonymous):

same.

OpenStudy (anonymous):

TBH, the first 5 attempts or so I had the original problem written wrong, that's how bad it is getting.

myininaya (myininaya):

wait

OpenStudy (anonymous):

Oh wait. I have it.

myininaya (myininaya):

ok polpak i think i got it we need to use partial fractions

OpenStudy (anonymous):

I'll look at yours, but I did mine much easier once I had the problem written right.

OpenStudy (anonymous):

It's looking pretty good! Check this out though: \[sinh(x) = \frac{e^x - e^{-x}}{2}\]\[cosh(x) = \frac{e^x + e^{-x}}{2}\] \[\implies \int \frac{e^{-x}- e^x}{e^{2x} + e^{-2x} + 2}dx\]\[=\int\frac{-2sinh(x)}{(e^x + e^{-x})^2}dx\]\[=\int\frac{-2sinh(x)}{[2cosh(x)]^2}dx\]\[=-\frac{1}{2}\int\frac{sinh(x)}{cosh^2(x)}dx\]Let \(u = cosh(x) \implies du = sinh(x)dx\)\[\implies -\frac{1}{2}\int\frac{sinh(x)}{cosh^2(x)}dx\]\[=-\frac{1}{2}\int\frac{1}{u^2}du = \frac{1}{2}sech(x) +C\]

OpenStudy (anonymous):

Yay we got the same answer!

myininaya (myininaya):

:)

OpenStudy (anonymous):

Nice job!

myininaya (myininaya):

polpak why did we do this

myininaya (myininaya):

im so tired

OpenStudy (anonymous):

Because after a certain point you're too tired to make good decisions.

myininaya (myininaya):

lol

OpenStudy (anonymous):

And this problem came up at just the critical time where we had made a good decision, then saw this a moment later.

OpenStudy (anonymous):

once we were no longer capable.

OpenStudy (anonymous):

Night Myin.

myininaya (myininaya):

night

OpenStudy (anonymous):

I mean it this time! Go to sleep ;p

myininaya (myininaya):

are you sure we got the same answer i can't check this how did you they were the same

myininaya (myininaya):

i mean i can check it but i cant right now

OpenStudy (anonymous):

Divide top and bottom of yours by e^x